OpenStudy (anonymous):

After a discount of 10%, the price of a jacket, including tax, is $70.79. The following equation can be used to find x, the price of the jacket without the discount. x - 0.1x + 0.06(x - 0.1x) = 70.79 What is the price of the jacket without the discount?

OpenStudy (anonymous):

74.2$ as x-0.1x(1+0.06)=70.79 x(1-0.1)=66.78 x=66.78/0.9 x=74.2$